### v1

# RoboHanger: Learning Generalizable Robotic Hanger Insertion for Diverse Garments

Abstract
For the task of hanging clothes, learning how to insert a hanger into a garment is crucial but has been seldom explored in robotics. In this work, we address the problem of inserting a hanger into various unseen garments that are initially laid out flat on a table. This task is challenging due to its long-horizon nature, the high degrees of freedom of the garments, and the lack of data. To simplify the learning process, we first propose breaking the task into several stages. Then, we formulate each stage as a policy learning problem and propose low-dimensional action parameterization. To overcome the challenge of limited data, we build our own simulator and create 144 synthetic clothing assets to effectively collect high-quality training data. Our approach uses single-view depth images and object masks as input, which mitigates the Sim2Real appearance gap and achieves high generalization capabilities for new garments. Extensive experiments in both simulation and the real world validate our proposed method. By training on various garments in the simulator, our method achieves a 75% success rate with 8 different unseen garments in the real world.
I Introduction
Laundry is a common and routine household chore that is time-consuming. The entire process involves many steps, including collecting clothes, washing, drying, hanging, unfolding, and folding. Therefore, automating laundry tasks with robots is highly appealing. Researchers have explored how to enable robots to fold and unfold clothes from arbitrary initial configurations [1, 2, 3], or to hang rigid or deformable objects on racks or hooks [4, 5, 6, 7]. However, we have identified that one crucial step in the laundry process — inserting a hanger into clothes — has been rarely explored by researchers.
In this work, we investigate how to use a dual-arm robot to insert a hanger into various unseen garments, as illustrated in Figure LABEL:fig:teaser. Our task begins with the target garment laid flat on the table. The goal is to insert the hanger into the garment and hang it up in the final state. End-to-end imitation learning methods face significant challenges in achieving satisfactory results due to (1) task complexity: inserting a hanger into a garment involves intricate, highly coordinated dual-arm actions in a long horizontal manner; (2) data scarcity: collecting large amounts of real-world demonstrations is inefficient and potentially dangerous, while existing simulators lack the necessary physical accuracy and diversity in suitable garments and hangers; and (3) generalizability: garments vary widely in color, shape, size, and initial configurations, making it difficult for models to generalize across different scenarios.
To address the three aforementioned issues, we propose a method that breaks the task into several stages. An overview of the entire process is provided in Figure 1. For different stages, we leverage both learnable and non-learnable modules to take the best of both worlds. Learnable modules offer more flexible and adaptive actions, while non-learnable modules reduce learning difficulty and improve robustness. We design neural networks to learn actions that need to adapt to new clothing shapes and initial configurations, including pressing, lifting, dragging, and rotating, as detailed in Section III-C. For actions that are less critical to the task’s success, we use fixed motions, such as insertion and pulling actions.
We build a simulator and create over 1,000 hanger and 100 garment assets for data collection. Our simulation environment is built upon Taichi [8], which supports large-scale parallel simulations on GPU. We choose depth and the object’s mask as input instead of raw RGB data to mitigate generalization issues. Depth images provide the necessary geometric information for the policies and are naturally unaffected by the varying colors and patterns of different garments. Extensive experiments in the simulation and real world validate our proposed method. Trained on 120 different garments in the simulation, our method can achieve a final success rate of 75% on 8 different unknown garments in the real world without any fine-tuning with real world data.
II Related Work
II-A Deformable Object Manipulation
Deformable object manipulation is a long-standing challenging topic due to the infinite object states and complex dynamics. Some comprehensive surveys can be found in [9, 10, 11]. One type of deformable object manipulation task involves handling one-dimensional objects such as ropes and cables [12, 13, 14, 15, 16, 17]. 2D clothing-like object manipulation tasks include cloth smoothing [18, 19, 20, 1, 21], cloth folding [22, 23, 2], cloth canonicalization [24, 3], bedding manipulation [25, 26] and bag [27, 28, 29] manipulation. Other works study 3D deformable object manipulation, such as solid object [30, 31] manipulation.
From the aspect of the method, early studies often use the model-based method for planning [32, 33, 34]. These kinds of methods often suffer from inaccurate dynamic models. Some methods have also proposed learning the dynamics model or representations of the deformable objects [35, 36, 37, 38, 39]. Recently, more and more research has explored data-driven approaches for learning policies such as imitation learning [40, 41, 42, 43] and reinforcement learning [6, 44, 45, 46]. Some researchers have also used differentiable simulators for deformable object manipulation tasks [47, 48]. In this work, we employ a data-driven approach, utilizing self-supervised trials to train our value network.
II-B Garment Manipulation
Garment manipulation is an important topic in deformable object manipulation. Folding [49, 50, 2] and unfolding [51, 52, 53, 54, 55] are two basic and widely studied directions. Many works also investigated robotic assisted dressing [56, 57, 58, 59]. Hanging garments is an extremely under-explored area. Some related works include: [6] studied using reinforcement learning to hang a fabric on a rack, [7] learned to hang a knit cap on a coat rack, [60] used a rule based method to manipulate the hanger. In this work, we use learning-based method to enable the robot to insert the hanger into the clothes through the neckline.
III Method
We focus on the task of inserting a hanger into a garment. As illustrated in Figure LABEL:fig:teaser, the initial phase involves a hanger and a garment placed separately on the table in front of the robot. The goal is to control a dual-arm mobile robot that inserts the hanger into the garment through the neckline. The final phase involves the robot lifting the hanger with the garment successfully hanging. To focus on the hanger insertion task, we implement simple heuristic policies for the first and final phases, especially, placing the garment faced up to allow its neckline observable. We also omit the control of the robot’s base, as it is beyond the scope of this work. The robot is manually positioned at the table and remains stationary during execution. Each of the robot’s 7-DoF arms is equipped with parallel grippers, and a depth camera is mounted on its head for perception, as shown in Figure III.
Figure 0: Robot system in the real world. The dual-arm robot is equipped with parallel grippers on both arms and uses a head-mounted camera for observation. The robot’s waist can be controlled to lean forward.
Figure 1: System Overview. Our system takes RGB-D observations. First, we segment the hanger, the garment, and its neckline, and detect the three keypoints of the hanger. The depth map, along with the masks, is then fed into our action primitive networks $Q_{1-4}$. Each network predicts a 2D keypoint $P_{1-4}$ in the image, indicating where the action primitives should operate. These keypoints are first back-projected into 3D space and then sent to the whole-body motion controller to execute the long-horizontal hanger insertion task step-by-step.
We will illustrate the overall pipeline in Section III-A and describe each module including: hanger detection in Section III-B, hanger insertion policy in Section III-C, and action primitive networks in Section III-D. The simulation and data collection process is given in Section III-E and the whole body motion controller is introduced in Section III-F.
III-A System Overview
Our method takes RGB-D images as input, as illustrated in Figure 1. We first detect three keypoints of the hanger: a grasping point, and the left and right endpoints. The grasping point indicates where the gripper should grasp the hanger for subsequent motions, while the left and right endpoints mark the locations where the hanger should be inserted into the target garment. Next, we use GroundedSAM [61] to segment the hanger, garment, and its neckline. When the raw depth data is of low quality, we use [62] to obtain a refined depth map. The refined depth map, along with the masks, is then concatenated and fed into action primitive networks to predict two critical action primitives: press-and-lift and drag-and-rotate, which are used to insert the left and right keypoints of the hanger into the garment, respectively.
III-B Garment Segmentation and Hanger Detection
Given a frame of aligned color and depth images, we first use GroundedSAM [61] to segment and obtain the hanger’s mask. Then, using the camera’s intrinsic and extrinsic parameters, we also obtain the hanger’s point cloud as shown in Figure 1. We design a simple network consisting of a PointNet++ backbone and a 3-layer MLP detection head which outputs three keypoints $\{H_{g},H_{l},H_{r}\}$. $H_{g}$ is the grasping point and $H_{l},H_{r}$ denotes the left and right endpoint of the hanger respectively. To train this network, We synthesize 1,000 meshes of hangers with annotated keypoints.
III-C Hanger Insertion Policy
Action primitives [1, 28, 63] are parameterized end-effector motions specifically designed to accomplish robotic tasks. They are effective for learning and robust in execution. Our hanger insertion policy consists of two key action primitives: press-and-lift and drag-and-rotate, each responsible for inserting the left and right ends of the hanger into the garment. Each action primitive is parameterized by two keypoints. Such action primitives are carefully designed to simplify learning and improve the success rate.
Press-and-lift
As illustrated in Figure 2, after successfully grasping the hanger, we first use the left endpoint $H_{l}$ of the hanger to press the neckline at point $P_{1}$. With the garment pressed by the hanger, we then lift the garment at point $P_{2}$ with the other hand without moving the garment. Finally, we insert the hanger into the garment to a fixed distance. By combining the ’press’ and ’lift’ actions, we can overcome the challenge of the gripper grasping both layers of the garment, which may hinder the insertion of the hanger.
Figure 2: Action primitive: press-and-lift. After grasping the hanger with the right hand, we first press the garment at point $P_{1}$ with the hanger, then lift the garment at point $P_{2}$, and finally insert the hanger to a fixed distance.
Drag-and-rotate
In the second action primitive, as illustrated in Figure 3, we first move the left gripper to the position where the right gripper previously released the hanger and re-grasp the hanger to avoid self-collision between the two arms. Next, we use the right gripper to grasp the clothing at point $P_{3}$ and drag it forward to the right. Then, we use the left gripper to rotate the hanger while moving the right end to $P_{4}$ and aligning the angle. Finally, in real-world experiments, we slightly pull the clothing back to arrange it properly.
Figure 3: Action primitive: drag-and-lift. We first switch to the left hand to re-grasp the hanger, avoiding self-collision during subsequent actions. Next, we use the right hand to drag the garment and rotate the hanger into it. Finally, we pull the garment to arrange it properly.
Success criteria and reward
For each action primitive, the success criterion is whether the hanger’s endpoint ($H_{l}$ for the first action primitive and $H_{r}$ for the second) is inserted into the garment. Similarly, if both endpoints are inserted, the task is considered successfully completed. This success criterion also serves as a sparse reward for training our action primitive networks. In simulation, we cast multiple rays randomly from the endpoint in all directions. If the proportion of rays intersecting with the garment’s mesh exceeds a threshold (0.95 for the left endpoint and 0.9 for the right in our experiments), we consider the endpoint to be covered by the garment.
III-D Action Primitive Networks
We use separate value networks to learn the parameters (operation points) for the two action primitives. Specifically, each network learns an action value map using a UNet architecture [64], as illustrated in Figure 1. Each pixel in this map corresponds to the action value of operating on that point. The network input, denoted as $O$, consists of a top-down depth map and object masks. Specifically, we use the garment mask and neckline mask when predicting $P_{1}$ and $P_{2}$ for the press-and-lift action, and the garment mask and hanger mask when predicting $P_{3}$ and $P_{4}$ for the drag-and-rotate action.
To train the four value networks, we simulate over 10,000 trajectories and their corresponding ground truth rewards. Each data entry is of the form $(O,P_{1},P_{2},R)$ where $P_{1}=(X_{1},Y_{1})$ and $P_{2}=(X_{2},Y_{2})$ correspond to each pixel in the output value map, and $R$ is the sparse reward received after executing these actions. The second value network $Q_{2}$ can be directly supervised using the sparse reward $R$ at the end of the execution of both key steps:
$\mathcal{L}_{2}=BCE(Q_{2}[P_{2};O,P_{1}],R)$ (1)
where $BCE$ denotes the binary cross-entropy loss and []delimited-[][$\cdot$][ ⋅ ] is the indexing operation. The input to the $Q_{2}$ is the observation $O$ concatenated with the one-hot map of $P_{1}$. The $Q_{2}$ network serves two purposes here: first, it can predict the value map for $P_{2}$ given $P_{1}$; second, the maximum of its output indicates the action value of point $P_{1}$. Therefore, we propose to use the maximum of the second network’s output $\hat{R}=(\max\{Q_{2}[\cdot;O,P_{1}]\}>0.5)$ as pseudo label for $Q_{1}$. Here $Q_{2}$ is the output after the Sigmoid function. The final loss function for the first network $Q_{1}$ is:
$\mathcal{L}_{1}=BCE(Q_{1}[P_{1};O],\hat{R})$ (2)
We apply a similar data collection and training strategy to train the value networks $Q_{3}$ and $Q_{4}$. When executing these action primitives, we back-project pixels into three-dimensional coordinates using depth value.
III-E Data Collection
We observe that existing simulators, such as PyFlex [65] and MuJoCo [66], either lack the accuracy or efficiency needed for our task. Their main issue is that the self-collisions in the clothing is not well handled. Therefore, we build our simulation environment based on Taichi [8].
Simulation environment and assets
We use FEM and implicit time integration with a single-step Newton iteration [67] for physical simulation. Cloth self-collision and cloth-to-rigid-body collision are handled with a cubic energy function. Cloth self-intersection is prevented using Continuous Collision Detection (CCD). The collision force between cloth and rigid bodies is computed through the Signed Distance Function (SDF) of the latter. We do not perform dynamics simulation or collision handling for rigid bodies and articulated bodies; instead, we compute the object’s position and velocity using kinematic relationships. We render depth images with SAPIEN [68] and design the 3D models of clothes and hangers (without texture) heuristically.
Domain randomization
We randomize the Young’s modulus and density of the clothing, as well as the friction coefficient of the garment in the simulation. Additionally, we randomize the camera’s intrinsic and extrinsic parameters to compensate for calibration errors when the system is deployed in the real world.
Data collection strategy
We use an iterative data collection approach. Specifically, we first design a state-based heuristic policy to collect trajectories, and use this data to train our networks for initialization. The heuristic policy attempts to find an action that meets all of our predefined rules through sampling. These rules are based on geometric information of the current clothing mesh, such as the number of layers of the garment near the grasping point. Next, we use the network policy to collect more data in the simulator, incorporate this online data into the dataset, and iteratively finetune the network.
III-F Whole-body Motion Controller
Since our hanger insertion task requires a large working space — with some actions necessitating the agent to lean forward without colliding with the table — we implement a custom whole-body controller to ensure the action space does not become a bottleneck. Specifically, after specifying the position of the robotic arm’s end effector, we only constrain the grippers’ fingertips to remain parallel to the table, leaving the other two rotational degrees of freedom unrestricted. Additionally, for steps that require an even larger workspace, we adjust the waist joint angles to allow the robot to lean forward. These angles are computed using a linear function based on the target positions of the left and right end effectors.
IV Experiment
IV-A Experiment Setup
Hanger. We find that using parallel grippers makes it difficult to reliably grasp thin hangers. Therefore, we employ custom-designed 3D-printed hangers to facilitate grasping. The primary difference between our hangers and regular ones is that our design includes a section specifically made for secure gripping.
Metrics. We use the success rate of inserting the hanger, $S$, to evaluate the performance of the policy. $S$ can be further divided into: the success rate of inserting the left endpoint, $S_{1}$; and the success rate of inserting the right endpoint after successfully inserting the left endpoint, $S_{2}$. In our policy, the left end is always inserted first, followed by the right end. Therefore, $S=S_{1}\times S_{2}$.
[…]
IV-B Comparison with Baselines in Simulation
We first compare our methods with several other baselines in simulation. Among the 144 synthetic clothing meshes, we randomly divided them into 120 training garments and 24 testing garments. The hangers used for training and testing are the ones shown in Figure 4 (a.1) and (a.2).
Methods Input Output S1(%) $\uparrow$ S2(%) $\uparrow$ S(%) $\uparrow$
Ours V Prim $88.7$ $\mathbf{95.9}$ $\mathbf{85.1}$
Heuristic S Prim $59.1$ $75.4$ $44.5$
IL V Prim $85.8$ $91.7$ $78.7$
SAC [69] V Prim $37.9$ $56.0$ $21.2$
ACT [43] V Joint $\mathbf{94.8}$ $67.0$ $63.5$
DP [70] V Joint $82.8$ $59.7$ $49.5$
TABLE I: Comparison with baseline policies in simulation.
We present the comparisons in Table I. ”Input” refers to the type of input used by the policy, which can be either visual input (V) or state input (S). ”Output” refers to the type of output generated by the policy, which can either be action primitives (Prim) or joint angles (Joint).
”Heuristic” represents a heuristic algorithm that we manually designed based on the current mesh of the clothing.
”IL” stands for imitation learning, which is also based on action primitives and UNets. We treat the task as a multi-classification task (selecting a pixel in the image) using cross-entropy loss and only retain positive samples for training the UNets.
”SAC” stands for Soft Actor-Critic algorithm [69]. In SAC, the input to the value network consists of both the observation and the action. The action is first transformed into an image where each pixel value represents the inverse of the distance to that action. This action image is then concatenated with the observation and fed into a CNN. The action network is also a CNN that directly outputs $(X_{i},Y_{i})$. Compared to our method, SAC requires learning an additional action network, while our method directly finds the maximum position on the output of the value network.
”ACT” and ”DP” (diffusion policy) are methods proposed by [43] and [70]. We directly use data generated by heuristics to train an end-to-end imitation learning policy. The network’s input consists of the raw depth images from the head camera as well as the masks of the clothing and the hanger.
Our method performs best in simulation and achieves an overall 85.1% success rate. ACT performs excellently when inserting the first endpoint, but due to accumulated errors, it performs poorly when inserting the second endpoint. Similar issues are observed in other end-to-end imitation learning baselines such as diffusion policy. We also observe that SAC performs worst probably due to the additional action network not sufficiently trained.
IV-C Ablation Studies in Simulation
We provide the ablation results in Table II to validate our design choices.
S1(%) $\uparrow$ S2(%) $\uparrow$ S(%) $\uparrow$
Ours $\mathbf{88.7}$ $\mathbf{95.9}$ $\mathbf{85.1}$
w/o online data $86.5$ $94.7$ $81.8$
w/o neckline mask $85.6$ $94.2$ $80.6$
TABLE II: Ablation study in simulation.
”w/o online data” indicates training the network using only offline data without online data, under the same data quantity. Fine-tuning with online data leads to an improvement in the final success rate. ”w/o neckline mask” indicates training the policy without using the neckline mask. It is evident that the neckline mask also contributes to an increase in the final success rate.
IV-D Comparison with Baselines in Real World
In real-world experiments, we test 8 different garments and Hanger 1, as shown in Figure 4. For a fair comparison, we use the robot to initialize the garment configuration. Additionally, to evaluate our policy’s performance under different initializations, we also report the policy’s success rate with human initialization.
Output S1(%) $\uparrow$ S2(%) $\uparrow$ S(%) $\uparrow$
Ours Prim $\mathbf{93.8}$ $80.0$ $\mathbf{75.0}$
IL Prim $75.0$ $\mathbf{83.3}$ $62.5$
ACT Joint $12.5$ $0.0$ $0.0$
Fix Prim $75.0$ $41.7$ $31.2$
  Ours(H) Prim $93.8$ $85.3$ $80.0$
TABLE III: Experiments in the real world.
We present the performance of our policy compared to other methods on the real robot in Table III. We include ”IL” (imitation learning), which performed best among baseline policies in simulation; ”ACT” (a policy that directly outputs joint angles); and ”Fix” (a policy outputs hard coded parameters of both action primitives) as baselines. Our method outperforms all baseline methods in terms of the final success rate which demonstrates the efficacy of our method.
Although the ACT policy shows a certain level of success in simulation, with a particularly high success rate when inserting the first endpoint, it suffers from a severe sim-to-real gap and almost never succeeds in the real world.
Fixed policy is a manually tuned policy for the test garments. It can be observed that the fixed policy still shows high success rates for inserting the left endpoint because the initial position of the clothes is relatively fixed in our experiments. However, when inserting the right endpoint, the success rate of the fixed policy drops significantly due to unforeseen deformations of the clothes during the left endpoint insertion.
In the last row of the table (”Ours(H)”), we further test the performance of our policy when the garment is initialized by a human. When the garment is initialized by the robot, the success rate shows only a slight decrease compared to human initialization, demonstrating the robustness of our policy to the initialization process.
IV-E Generalization of Hangers in Real World
To further study the generalizability of our method in terms of different hangers, we report the performance of our policy when using different hangers in Table IV. Here, Hanger 2 is a hanger seen during training but without a crossbar. Hanger 3 and Hanger 1 have the same geometry, but due to difference settings in the 3D printing process, Hanger 3’s surface is much rougher. Hanger 4 is a hanger that is not seen during training.
S1(%) $\uparrow$ S2(%) $\uparrow$ S(%) $\uparrow$
Hanger 1 $\mathbf{93.8}$ $\mathbf{80.0}$ $\mathbf{75.0}$
Hanger 2 $81.2$ $23.1$ $18.8$
Hanger 3 $75.0$ $75.0$ $56.2$
Hanger 4 $\mathbf{93.8}$ $66.7$ $62.5$
TABLE IV: Comparison of performance with different hangers in the real world.
Although Hanger 2 was seen during training, there is a significant performance drop when inserting the right endpoint. As shown in Figures 5(a) and 5(b), these two images compare the state when inserting the right endpoint using different hangers. The presence of a crossbar helps the hanger separate the upper and lower layers of the clothing, making insertion easier. When using Hanger 3, as shown in Figures 5(c) and 5(d), if the hanger has too much friction, it may drag the clothes while moving, causing the clothes to become very wrinkled, which in turn reduces the success rate. When using Hanger 4, the complexity of the hanger’s shape also affects the final success rate.
Figure 5: Results with different hangers in the real world. In the real world, issues may arise when changing hangers. Figure (a) and Figure (c) show results with Hanger 1. As shown in Figure (b), when switching to Hanger 2, it becomes difficult to separate the upper and lower layers of the clothing while inserting the right end. Figure (d) illustrates that with Hanger 3, which has higher friction, the clothing tends to get wrinkled during the insertion of the left end, causing the garment to become stuck in a worse state.
V Limitation
For garments with less elasticity, such as shirts, more precise force feedback control might be required to avoid damage to the hanger and garment. Another challenge could arise when applying our action primitives to irregularly shaped garments, such as a shirt with a hood.
VI Conclusion
In this work, we investigate the task of inserting a hanger into the neckline of clothes. For such a challenging but rarely studied task, we propose a simulation environment and data collection pipeline to address it. We demonstrate the efficiency of training the algorithm in a simulated environment, as well as its effectiveness and generalizability in the real world. Our pipeline facilitates efficient demonstration collecting for future robotic foundation models that require large amounts of diverse data. Future work may involve relaxing the assumptions for this task and extending the method to handle a wider variety of garments.

### v4

@@
 Abstract
-For the task of hanging clothes, learning how to insert a hanger into a garment is crucial but has been seldom explored in robotics.
-In this work, we address the problem of inserting a hanger into various unseen garments that are initially laid out flat on a table.
-This task is challenging due to its long-horizon nature, the high degrees of freedom of the garments, and the lack of data.
-To simplify the learning process, we first propose breaking the task into several stages.
-Then, we formulate each stage as a policy learning problem and propose low-dimensional action parameterization.
+For the task of hanging clothes, learning how to insert a hanger into a garment is a crucial step, but has rarely been explored in robotics.
+In this work, we address the problem of inserting a hanger into various unseen garments that are initially laid flat on a table.
+This task is challenging due to its long-horizon nature, the high degrees of freedom of the garments and the lack of data.
+To simplify the learning process, we first propose breaking the task into several subtasks.
+Then, we formulate each subtask as a policy learning problem and propose a low-dimensional action parameterization.
 To overcome the challenge of limited data, we build our own simulator and create 144 synthetic clothing assets to effectively collect high-quality training data.
 Our approach uses single-view depth images and object masks as input, which mitigates the Sim2Real appearance gap and achieves high generalization capabilities for new garments.
-Extensive experiments in both simulation and the real world validate our proposed method.
+Extensive experiments in both simulation and reality validate our proposed method.
 By training on various garments in the simulator, our method achieves a 75% success rate with 8 different unseen garments in the real world.
 I Introduction
-Laundry is a common and routine household chore that is time-consuming.
-The entire process involves many steps, including collecting clothes, washing, drying, hanging, unfolding, and folding.
-Therefore, automating laundry tasks with robots is highly appealing.
-Researchers have explored how to enable robots to fold and unfold clothes from arbitrary initial configurations [1, 2, 3], or to hang rigid or deformable objects on racks or hooks [4, 5, 6, 7].
-However, we have identified that one crucial step in the laundry process — inserting a hanger into clothes — has been rarely explored by researchers.
-In this work, we investigate how to use a dual-arm robot to insert a hanger into various unseen garments, as illustrated in Figure LABEL:fig:teaser.
-Our task begins with the target garment laid flat on the table.
-The goal is to insert the hanger into the garment and hang it up in the final state.
-End-to-end imitation learning methods face significant challenges in achieving satisfactory results due to (1) task complexity: inserting a hanger into a garment involves intricate, highly coordinated dual-arm actions in a long horizontal manner; (2) data scarcity: collecting large amounts of real-world demonstrations is inefficient and potentially dangerous, while existing simulators lack the necessary physical accuracy and diversity in suitable garments and hangers; and (3) generalizability: garments vary widely in color, shape, size, and initial configurations, making it difficult for models to generalize across different scenarios.
-To address the three aforementioned issues, we propose a method that breaks the task into several stages.
-An overview of the entire process is provided in Figure 1.
-For different stages, we leverage both learnable and non-learnable modules to take the best of both worlds.
-Learnable modules offer more flexible and adaptive actions, while non-learnable modules reduce learning difficulty and improve robustness.
-We design neural networks to learn actions that need to adapt to new clothing shapes and initial configurations, including pressing, lifting, dragging, and rotating, as detailed in Section III-C.
-For actions that are less critical to the task’s success, we use fixed motions, such as insertion and pulling actions.
-We build a simulator and create over 1,000 hanger and 100 garment assets for data collection.
-Our simulation environment is built upon Taichi [8], which supports large-scale parallel simulations on GPU.
+As a common household task, using a hanger to hang clothes is a challenging deformable manipulation problem for robots.
+It involves complex contact dynamics between the hanger and the clothes and requires precise coordination between two robotic arms.
+The entire task consists of flattening the clothes, inserting the hanger, and hanging the hanger on a rack.
+Researchers have explored methods to flatten clothes from arbitrary initial configurations [1, 2], as well as how to hang objects on racks or hooks [3, 4, 5].
+However, we have identified that one crucial step—inserting the hanger into the garment—has rarely been studied.
+In recent years, learning-based approaches for deformable object manipulation tasks have made significant progress.
+Zhao et al.
+[6] employ end-to-end imitation learning and collect data through teleoperation in the real world to accomplish this task.
+However, this approach requires substantial human labor for data collection during the training process.
+In this work, we investigate how to use a dual-arm robot to insert a hanger into various unseen garments, as illustrated in Figure LABEL:fig:teaser, employing a Sim2Real approach and without any human labeling effort.
+To focus on the hanger insertion process, the task begins with the target garment laid flat on the table, a scenario that has been addressed in several previous works [1, 2].
+This task faces three significant challenges: (1) task complexity: the action space is large and it is difficult to define a good reward function, which makes it hard to find a good policy through sampling by directly applying reinforcement learning algorithms; (2) data scarcity: existing simulators lack the necessary physical accuracy and diversity in terms of suitable garments and hangers; and (3) generalizability: garments vary widely in color, shape, and size, making generalization across different garments a non-trivial problem.
+To address the three aforementioned challenges, we propose a method that decomposes the task into two subtasks: inserting the left end of the hanger and inserting the right end of the hanger.
+Each subtask is accomplished using an action primitive [7, 8, 9], which is a simple trajectory parameterized by two 2D keypoints.
+These four 2D keypoints are learned through four UNet [10] models.
+This low-dimensional parameterization facilitates the learning process and ensures robustness during policy execution in both simulation and reality.
+To train the models, we build a simulator and generate over 1,000 hangers and 100 garments for data collection.
+Our simulation environment is built upon Taichi [11], which supports large-scale parallel simulations on GPUs.
+We generate training data through trial and error within the simulator, minimizing the need for human labeling effort.
 We choose depth and the object’s mask as input instead of raw RGB data to mitigate generalization issues.
 Depth images provide the necessary geometric information for the policies and are naturally unaffected by the varying colors and patterns of different garments.
-Extensive experiments in the simulation and real world validate our proposed method.
-Trained on 120 different garments in the simulation, our method can achieve a final success rate of 75% on 8 different unknown garments in the real world without any fine-tuning with real world data.
+Extensive experiments in simulation and the real world validate our proposed method.
+Trained on 120 different garments in simulation, our method can achieve a final success rate of 75% on 8 different unseen garments in reality without any fine-tuning with real data.
 II Related Work
-II-A Deformable Object Manipulation
-Deformable object manipulation is a long-standing challenging topic due to the infinite object states and complex dynamics.
-Some comprehensive surveys can be found in [9, 10, 11].
-One type of deformable object manipulation task involves handling one-dimensional objects such as ropes and cables [12, 13, 14, 15, 16, 17].
-2D clothing-like object manipulation tasks include cloth smoothing [18, 19, 20, 1, 21], cloth folding [22, 23, 2], cloth canonicalization [24, 3], bedding manipulation [25, 26] and bag [27, 28, 29] manipulation.
-Other works study 3D deformable object manipulation, such as solid object [30, 31] manipulation.
-From the aspect of the method, early studies often use the model-based method for planning [32, 33, 34].
-These kinds of methods often suffer from inaccurate dynamic models.
-Some methods have also proposed learning the dynamics model or representations of the deformable objects [35, 36, 37, 38, 39].
-Recently, more and more research has explored data-driven approaches for learning policies such as imitation learning [40, 41, 42, 43] and reinforcement learning [6, 44, 45, 46].
-Some researchers have also used differentiable simulators for deformable object manipulation tasks [47, 48].
-In this work, we employ a data-driven approach, utilizing self-supervised trials to train our value network.
-II-B Garment Manipulation
-Garment manipulation is an important topic in deformable object manipulation.
-Folding [49, 50, 2] and unfolding [51, 52, 53, 54, 55] are two basic and widely studied directions.
-Many works also investigated robotic assisted dressing [56, 57, 58, 59].
-Hanging garments is an extremely under-explored area.
-Some related works include: [6] studied using reinforcement learning to hang a fabric on a rack, [7] learned to hang a knit cap on a coat rack, [60] used a rule based method to manipulate the hanger.
-In this work, we use learning-based method to enable the robot to insert the hanger into the clothes through the neckline.
-III Method
-We focus on the task of inserting a hanger into a garment.
-As illustrated in Figure LABEL:fig:teaser, the initial phase involves a hanger and a garment placed separately on the table in front of the robot.
-The goal is to control a dual-arm mobile robot that inserts the hanger into the garment through the neckline.
-The final phase involves the robot lifting the hanger with the garment successfully hanging.
-To focus on the hanger insertion task, we implement simple heuristic policies for the first and final phases, especially, placing the garment faced up to allow its neckline observable.
-We also omit the control of the robot’s base, as it is beyond the scope of this work.
-The robot is manually positioned at the table and remains stationary during execution.
-Each of the robot’s 7-DoF arms is equipped with parallel grippers, and a depth camera is mounted on its head for perception, as shown in Figure III.
+II-A Learning-Based Cloth Manipulation
+Learning-based robotic cloth manipulation policies have recently gained wide interest.
+Many works collect data from human demonstrations [6, 9, 12, 1].
+The advantage of these methods is the absence of a domain gap, but they require a large amount of manual effort and often suffer from limited generalization.
+Sim-to-real transfer is a safe and efficient approach [2, 13, 14, 15].
+However, previous tasks typically focus only on manipulating the garment itself, without requiring the manipulation of objects other than the garment, such as folding [16, 17, 18] and unfolding [19, 2, 20].
+These tasks require a lower diversity of actions, making it easier for the robot to learn.
+In contrast, inserting a hanger into clothes is a long-horizon task that requires the manipulation of both hanger and garment: first, the hanger must be grasped, then both endpoints of the hanger must be inserted into the clothing, and finally, the hanger must be lifted.
+To efficiently generate a large amount of high-quality training data, we design action primitives to reduce the action space.
+We build our own cloth simulator to automatically generate high-quality training data and labels through trial and error in simulation.
+II-B Robotic Hanging
+Robotic hanging is a task of significant practical importance, yet it has rarely been studied.
+Previous works have mainly focused on hanging objects on racks or hooks [5, 21, 22], without using hangers for clothing.
+In these tasks, typically, only contact between stationary rigid bodies and clothing is involved.
+However, in our task, the robot needs to simultaneously manipulate both the hanger and the clothing, increasing the complexity of the task.
+Researchers [23] have also tried using a rule-based method to manipulate the hanger, but the method they use lacks sufficient generalizability.
+Additionally, the manipulation policy requires the hanger and clothing to be pre-attached to the robotic arm’s end effector.
+In our task setup, both the garment and the hanger are initially placed on the table, and the grasping pose is output by our policy.
+Our manipulation policy for robotic hanger insertion demonstrates strong generalizability across different types of garments.
 Figure 0: Robot system in the real world.
@@
 The robot’s waist can be controlled to lean forward.
+III Method
 Figure 1: System Overview.
-Our system takes RGB-D observations.
-First, we segment the hanger, the garment, and its neckline, and detect the three keypoints of the hanger.
-The depth map, along with the masks, is then fed into our action primitive networks $Q_{1-4}$.
-Each network predicts a 2D keypoint $P_{1-4}$ in the image, indicating where the action primitives should operate.
-These keypoints are first back-projected into 3D space and then sent to the whole-body motion controller to execute the long-horizontal hanger insertion task step-by-step.
-We will illustrate the overall pipeline in Section III-A and describe each module including: hanger detection in Section III-B, hanger insertion policy in Section III-C, and action primitive networks in Section III-D.
-The simulation and data collection process is given in Section III-E and the whole body motion controller is introduced in Section III-F.
+(a) Before each action primitive, our system takes RGB-D observations as input and segments the hanger, the garment, and its neckline.
+We pre-detect three keypoints of the hanger at the beginning of the policy.
+(b) The action primitive press-and-lift inserts the left endpoint of the hanger into the target garment.
+(c) The action primitive drag-and-rotate inserts the right endpoint of the hanger into the target garment.
+In (b) and (c), four U-Nets ($Q_{press},Q_{lift},Q_{drag},Q_{rotate}$) all take the depth map and masks as input and output 2D value maps indicating the subtask success rate of applying actions at each pixel.
+We apply argmax to obtain a final single point.
+We focus on the task of inserting a hanger into a garment.
+As illustrated in Figure LABEL:fig:teaser, the initial state consists of a hanger and a garment placed separately on the table in front of the robot.
+The goal is to control a dual-arm robot to insert the hanger into the garment through the neckline.
+The final phase involves the robot lifting the hanger, with the garment successfully hanging.
+To focus specifically on the task of inserting the hanger, we use the method of previous work [2] to initialize the garment.
+After inserting the hanger into the clothes, we also omit the step of hanging the hanger on a rack.
+The robot is manually positioned in front of the table, and its base remains stationary during execution.
+Each of the 7-DoF arms of the robot is equipped with parallel grippers, and a depth camera is mounted on its head for perception, as shown in Figure II-B.
+We will illustrate the overall pipeline in Section III-A and describe each module, including hanger detection in Section III-B, hanger insertion policy in Section III-C, success criteria in Section III-D, action primitive networks in Section III-E, the simulation and data collection process in Section III-F, and the whole-body motion controller in Section III-G.
 III-A System Overview
 Our method takes RGB-D images as input, as illustrated in Figure 1.
-We first detect three keypoints of the hanger: a grasping point, and the left and right endpoints.
+We first detect three keypoints of the hanger: the grasping point $H_{g}$, the left endpoint $H_{l}$, and the right endpoint $H_{r}$.
 The grasping point indicates where the gripper should grasp the hanger for subsequent motions, while the left and right endpoints mark the locations where the hanger should be inserted into the target garment.
 Next, we use GroundedSAM [24] to segment the hanger, garment, and its neckline.
-When the raw depth data is of low quality, we use [62] to obtain a refined depth map.
 The refined depth map, along with the masks, is then concatenated and fed into action primitive networks to predict two critical action primitives: press-and-lift and drag-and-rotate, which are used to insert the left and right keypoints of the hanger into the garment, respectively.
 III-B Garment Segmentation and Hanger Detection
-Given a frame of aligned color and depth images, we first use GroundedSAM [61] to segment and obtain the hanger’s mask.
-Then, using the camera’s intrinsic and extrinsic parameters, we also obtain the hanger’s point cloud as shown in Figure 1.
-We design a simple network consisting of a PointNet++ backbone and a 3-layer MLP detection head which outputs three keypoints $\{H_{g},H_{l},H_{r}\}$.
-$H_{g}$ is the grasping point and $H_{l},H_{r}$ denotes the left and right endpoint of the hanger respectively.
+Given a frame of aligned color and depth images, we first use GroundedSAM [24] to segment and obtain the hanger mask.
+Using the intrinsic and extrinsic parameters of the camera, we can also obtain the point cloud of the hanger as shown in Figure 1(a).
+We design a simple network consisting of a PointNet++ backbone and a 3-layer MLP detection head that outputs three keypoints $\{H_{g},H_{l},H_{r}\}$.
 To train this network, we synthesize 1,000 meshes of hangers with annotated keypoints.
+Generating the data and training takes approximately 1 day to run on a single NVIDIA RTX 3090.
+After predicting these three keypoints, we grasp the hanger using a calculated grasping pose, with the position centered at the point $H_{g}$ and a rotation of $+45^{\circ}$ around the vector $(H_{l}-H_{r})$, as shown in the bottom left of Figure 1(b).
 III-C Hanger Insertion Policy
-Action primitives [1, 28, 63] are parameterized end-effector motions specifically designed to accomplish robotic tasks.
+Action primitives [7, 8, 9, 25] are parameterized end-effector motions specifically designed to accomplish complex robotic tasks.
 They are effective for learning and robust in execution.
-Our hanger insertion policy consists of two key action primitives: press-and-lift and drag-and-rotate, each responsible for inserting the left and right ends of the hanger into the garment.
-Each action primitive is parameterized by two keypoints.
-Such action primitives are carefully designed to simplify learning and improve the success rate.
+Our hanger insertion policy consists of two action primitives: press-and-lift and drag-and-rotate, each responsible for inserting the left and right ends of the hanger into the garment.
 press-and-lift
-As illustrated in Figure 2, after successfully grasping the hanger, we first use the left endpoint $H_{l}$ of the hanger to press the neckline at point $P_{1}$.
-With the garment pressed by the hanger, we then lift the garment at point $P_{2}$ with the other hand without moving the garment.
-Finally, we insert the hanger into the garment to a fixed distance.
-By combining the ’press’ and ’lift’ actions, we can overcome the challenge of the gripper grasping both layers of the garment, which may hinder the insertion of the hanger.
+This action primitive is parametrized by $(p_{press},p_{lift})=(x_{press},y_{press},x_{lift},y_{lift})\in\mathbb{R}^{4}$.
+As illustrated in Figure 2, after successfully grasping the hanger, we first use the left endpoint $H_{l}$ of the hanger to press the neckline at point $p_{press}$.
+With the garment pressed by the hanger, we then lift the garment at point $p_{lift}$ with the other hand without moving the garment.
+Finally, we insert the hanger into the garment at a fixed distance.
+By combining the press and lift actions, we can overcome the challenge of the gripper grasping both layers of the garment, which may otherwise hinder the insertion of the hanger.
 Figure 2: Action primitive: press-and-lift.
-After grasping the hanger with the right hand, we first press the garment at point $P_{1}$ with the hanger, then lift the garment at point $P_{2}$, and finally insert the hanger to a fixed distance.
 drag-and-rotate
-In the second action primitive, as illustrated in Figure 3, we first move the left gripper to the position where the right gripper previously released the hanger and re-grasp the hanger to avoid self-collision between the two arms.
-Next, we use the right gripper to grasp the clothing at point $P_{3}$ and drag it forward to the right.
-Then, we use the left gripper to rotate the hanger while moving the right end to $P_{4}$ and aligning the angle.
-Finally, in real-world experiments, we slightly pull the clothing back to arrange it properly.
-Figure 3: Action primitive: drag-and-lift.
-We first switch to the left hand to re-grasp the hanger, avoiding self-collision during subsequent actions.
-Next, we use the right hand to drag the garment and rotate the hanger into it.
-Finally, we pull the garment to arrange it properly.
-Success criteria and reward
+This action primitive is parametrized by $(p_{drag},p_{rotate})=(x_{drag},y_{drag},x_{rotate},y_{rotate})\in\mathbb{R}^{4}$.
+As illustrated in Figure 3, we first move the left gripper to the position where the right gripper previously released the hanger and re-grasp the hanger to avoid self-collision between the two arms.
+Next, we use the right gripper to grasp the garment at point $p_{drag}$ and drag it forward to the right.
+Then, we use the left gripper to rotate the hanger while moving the right end to $p_{rotate}$ and aligning it to the correct angle.
+Finally, we slightly pull the garment back to arrange it properly.
+Figure 3: Action primitive: drag-and-rotate.
+III-D Success Criteria
 For each action primitive, the success criterion is whether the hanger’s endpoint ($H_{l}$ for the first action primitive and $H_{r}$ for the second) is inserted into the garment.
-Similarly, if both endpoints are inserted, the task is considered successfully completed.
-This success criterion also serves as a sparse reward for training our action primitive networks.
-In simulation, we cast multiple rays randomly from the endpoint in all directions.
+Similarly, if both endpoints are inserted, the task is considered completed.
+We use $S_{1}$ and $S_{2}$ to represent the success status of each action primitive.
+$S_{1}=1$ indicates the successful insertion of the left endpoint, and $S_{2}=1$ indicates the successful insertion of the right endpoint.
+$S_{1}=0$ or $S_{2}=0$ indicates failure.
+To detect successful insertion during simulation, we cast multiple rays randomly from the endpoint in different directions.
 If the proportion of rays intersecting with the garment’s mesh exceeds a threshold (0.95 for the left endpoint and 0.9 for the right in our experiments), we consider the endpoint to be covered by the garment.
-III-D Action Primitive Networks
-We use separate value networks to learn the parameters (operation points) for the two action primitives.
-Specifically, each network learns an action value map using a UNet architecture [64], as illustrated in Figure 1.
-Each pixel in this map corresponds to the action value of operating on that point.
-The network input, denoted as $O$, consists of a top-down depth map and object masks.
-Specifically, we use the garment mask and neckline mask when predicting $P_{1}$ and $P_{2}$ for the press-and-lift action, and the garment mask and hanger mask when predicting $P_{3}$ and $P_{4}$ for the drag-and-rotate action.
-To train the four value networks, we simulate over 10,000 trajectories and their corresponding ground truth rewards.
-Each data entry is of the form $(O,P_{1},P_{2},R)$ where $P_{1}=(X_{1},Y_{1})$ and $P_{2}=(X_{2},Y_{2})$ correspond to each pixel in the output value map, and $R$ is the sparse reward received after executing these actions.
-The second value network $Q_{2}$ can be directly supervised using the sparse reward $R$ at the end of the execution of both key steps:
-$\mathcal{L}_{2}=BCE(Q_{2}[P_{2};O,P_{1}],R)$ (1)
-where $BCE$ denotes the binary cross-entropy loss and []delimited-[][$\cdot$][ ⋅ ] is the indexing operation.
-The input to the $Q_{2}$ is the observation $O$ concatenated with the one-hot map of $P_{1}$.
-The $Q_{2}$ network serves two purposes here: first, it can predict the value map for $P_{2}$ given $P_{1}$; second, the maximum of its output indicates the action value of point $P_{1}$.
-Therefore, we propose to use the maximum of the second network’s output $\hat{R}=(\max\{Q_{2}[\cdot;O,P_{1}]\}>0.5)$ as pseudo label for $Q_{1}$.
-Here $Q_{2}$ is the output after the Sigmoid function.
-The final loss function for the first network $Q_{1}$ is:
-$\mathcal{L}_{1}=BCE(Q_{1}[P_{1};O],\hat{R})$ (2)
-We apply a similar data collection and training strategy to train the value networks $Q_{3}$ and $Q_{4}$.
-When executing these action primitives, we back-project pixels into three-dimensional coordinates using depth value.
-III-E Data Collection
+This proportion is similar to the generalized winding number [26] in $\mathbb{R}^{3}$ space.
+The thresholds are set manually by inspecting many successful and unsuccessful trajectories and meshes in simulation.
+III-E Action Primitive Networks
+For each action primitive, we utilize two networks to predict the actions of the left and right hands, respectively.
+In the first action primitive, we use the network $Q_{press}$ to predict the press position $p_{press}$, and the network $Q_{lift}$ to predict the lift position $p_{lift}$.
+Likewise, for the second action primitive, we use network $Q_{drag}$ to predict the drag position $p_{drag}$, and network $Q_{rotate}$ to predict the rotate position $p_{rotate}$.
+These networks are similar to action-value functions in reinforcement learning, but we only consider the success rate within a single subtask.
+The network architecture is based on UNet [10].
+The network input, denoted as $O$, consists of a depth image and object masks.
+Specifically, we use the garment mask and neckline mask when predicting $p_{press}$ and $p_{lift}$ for the press-and-lift action, and the garment mask and hanger mask when predicting $Q_{drag}$ and $Q_{rotate}$ for the drag-and-rotate action.
+The network output is an image of the same size as the input, with each pixel representing the success rate of inserting one endpoint of the hanger, as mentioned in III-D.
+We employ the following dataflow to collaboratively learn two networks in an action primitive.
+During inference, $Q_{press}$ predicts $p_{press}$, and then $Q_{lift}$ predicts $p_{lift}$ conditioning on $p_{press}$.
+During training, we use trained $Q_{lift}$ to provide supervision for $Q_{press}$.
+Specifically, given data of the form $(O_{1},p_{press},p_{lift},S_{1})$, where $O\in\mathbb{R}^{C\times H\times W}$ is the observation, $p_{press},p_{lift}\in\mathbb{R}^{2}$, $S_{1}\in\{0,1\}$ denotes success or failure.
+We use $Q[y;x]\in\mathbb{R}$ to represent the input image $x$ to the UNet, and select the $y$-th pixel in the output image.
+The loss function of $Q_{lift}$ is:
+$\mathcal{L}_{lift}=BCE(Q_{lift}[p_{lift};O_{1},p_{press}],S_{1})$ (1)
+where $BCE$ denotes the binary cross-entropy loss.
+The input $O_{1}^{\prime}=(O_{1},p_{press})\in\mathbb{R}^{(C+1)\times H\times W}$ to the network $Q_{lift}$ is modified to be $O_{1}$ concatenated with the one-hot map of $p_{press}$, which is used to represent the conditioning on $p_{press}$.
+The maximum value of the $Q_{lift}$’s output can serve as supervision for $Q_{press}$:
+$\hat{S_{1}}=\max\{Q_{lift}[\cdot;O_{1},p_{press}]\}>0.5$ (2)
+$\mathcal{L}_{press}=BCE(Q_{press}[p_{press};O_{1}],\hat{S_{1}})$ (3)
+Similarly, given data of the form $(O_{2},p_{drag},p_{rotate},S_{2})$, the loss functions of $Q_{drag}$ and $Q_{rotate}$ are:
+$\mathcal{L}_{rotate}=BCE(Q_{rotate}[p_{rotate};O_{2},p_{drag}],S_{2})$ (4)
+$\hat{S_{2}}=\max\{Q_{rotate}[\cdot;O_{2},p_{drag}]\}>0.5$ (5)
+$\mathcal{L}_{drag}=BCE(Q_{drag}[p_{drag};O_{2}],\hat{S_{2}})$ (6)
+We apply argmax to the resulting 2D value map to obtain a single final pixel.
+When executing these action primitives, we back-project pixels into the world frame.
+The parameters of the action primitives only use the x and y coordinates, while the z coordinate is fixed.
+This is to avoid failures caused by depth camera inaccuracies.
+III-F Data Collection
+Simulation environment and assets
 We observe that existing simulators, such as PyFlex [27] and MuJoCo [28], either lack the accuracy or efficiency needed for our task.
-Their main issue is that the self-collisions in the clothing is not well handled.
+The main physics accuracy issue is that they do not handle collisions and dynamics of clothing very well.
 Therefore, we build our simulation environment based on Taichi [11].
-Simulation environment and assets
-We use FEM and implicit time integration with a single-step Newton iteration [67] for physical simulation.
+We use the finite element method (FEM) and implicit time integration with a single-step Newton iteration [29] for the physical simulation.
 Cloth self-collision and cloth-to-rigid-body collision are handled with a cubic energy function.
 Cloth self-intersection is prevented using Continuous Collision Detection (CCD).
-The collision force between cloth and rigid bodies is computed through the Signed Distance Function (SDF) of the latter.
-We do not perform dynamics simulation or collision handling for rigid bodies and articulated bodies; instead, we compute the object’s position and velocity using kinematic relationships.
-We render depth images with SAPIEN [68] and design the 3D models of clothes and hangers (without texture) heuristically.
-Domain randomization
-We randomize the Young’s modulus and density of the clothing, as well as the friction coefficient of the garment in the simulation.
-Additionally, we randomize the camera’s intrinsic and extrinsic parameters to compensate for calibration errors when the system is deployed in the real world.
+The collision force between cloth and rigid bodies is computed using the Signed Distance Function (SDF) of the latter.
+We do not perform simulation of dynamics or collision handling for rigid and articulated bodies; instead, we compute the position and velocity of objects using kinematic relationships.
+Depth images are rendered with SAPIEN [30], and we design the 3D models of clothes and hangers heuristically.
 Data collection strategy
 We use an iterative data collection approach.
-Specifically, we first design a state-based heuristic policy to collect trajectories, and use this data to train our networks for initialization.
-The heuristic policy attempts to find an action that meets all of our predefined rules through sampling.
-These rules are based on geometric information of the current clothing mesh, such as the number of layers of the garment near the grasping point.
-Next, we use the network policy to collect more data in the simulator, incorporate this online data into the dataset, and iteratively finetune the network.
-III-F Whole-body Motion Controller
+Specifically, we first design a state-based heuristic policy to collect trajectories and use this dataset to initialize our networks.
+This policy achieves a success rate of approximately 50%, generating a balanced dataset with both positive and negative samples.
+The heuristic policy attempts to find actions that satisfy all of our predefined rules through sampling.
+Next, we use the network policy to collect more data in simulation, incorporate this online data into the dataset, and iteratively fine-tune the network.
+During the initialization phase, we collect 12,000 trajectories.
+In the subsequent fine-tuning phase, we collect 1,200 trajectories per iteration, with a total of 5 iterations, resulting in 18,000 trajectories in total.
+Simulation and rendering take approximately 360 GPU hours on an NVIDIA RTX 3090, and training the network requires around 1 hour.
+III-G Whole-body Motion Controller
 Since our hanger insertion task requires a large working space — with some actions necessitating the agent to lean forward without colliding with the table — we implement a custom whole-body controller to ensure the action space does not become a bottleneck.
@@
 These angles are computed using a linear function based on the target positions of the left and right end effectors.
+III-H Sim-to-real Transfer
+For sim-to-real visual gap, we randomize the camera’s intrinsic and extrinsic parameters, depth maps, and masks during network training.
+This noise enables the network to handle a wider range of input distributions.
+When deploying in the real world, we use the method from [31] to obtain a refined depth map.
+For sim-to-real physics gap, we randomize the Young’s modulus and the density of the clothing, as well as the coefficient of friction of the garment in the simulation.
+The head camera is carefully calibrated. Apart from this, we do not perform any additional real fine-tuning and directly transfer the policy learned in simulation to the real world.
 IV Experiment
-IV-A Experiment Setup
-Hanger.
-We find that using parallel grippers makes it difficult to reliably grasp thin hangers.
-Therefore, we employ custom-designed 3D-printed hangers to facilitate grasping.
-The primary difference between our hangers and regular ones is that our design includes a section specifically made for secure gripping.
-Metrics.
-We use the success rate of inserting the hanger, $S$, to evaluate the performance of the policy.
-$S$ can be further divided into: the success rate of inserting the left endpoint, $S_{1}$; and the success rate of inserting the right endpoint after successfully inserting the left endpoint, $S_{2}$.
+IV-A Metrics
+We evaluate the performance of the policy based on the success rate $S$ of hanger insertion.
+$S$ can be further divided into: the success rate of inserting the left endpoint, $S_{1}$, and the success rate of inserting the right endpoint after successfully inserting the left endpoint, $S_{2}$.
 In our policy, the left end is always inserted first, followed by the right end.
 Therefore, $S=S_{1}\times S_{2}$.
+The success rate is computed based on 720 trials in simulation and 16 trials in the real world.
 […]
 IV-B Comparison with Baselines in Simulation
-We first compare our methods with several other baselines in simulation.
+We first compare our method with several state-of-the-art reinforcement learning and imitation learning approaches in simulation.
 Among the 144 synthetic clothing meshes, we randomly divided them into 120 training garments and 24 testing garments.
@@
 We present the comparisons in Table I.
-”Input” refers to the type of input used by the policy, which can be either visual input (V) or state input (S).
-”Output” refers to the type of output generated by the policy, which can either be action primitives (Prim) or joint angles (Joint).
-”Heuristic” represents a heuristic algorithm that we manually designed based on the current mesh of the clothing.
-”IL” stands for imitation learning, which is also based on action primitives and UNets.
-We treat the task as a multi-classification task (selecting a pixel in the image) using cross-entropy loss and only retain positive samples for training the UNets.
-”SAC” stands for Soft Actor-Critic algorithm [69].
+Input refers to the type of input used by the policy, which can be either visual input (V) or state input (S).
+Output refers to the type of output generated by the policy, which can either be action primitives (Prim) or joint angles (Joint).
+Our method performs the best in simulation, achieving an overall success rate of 85.1%.
+Heuristic refers to a heuristic algorithm that we manually designed based on the current mesh of the garment, as described in III-F (c).
+Since these rules are manually defined, their performance is not as good as methods that learn good actions from large amounts of data based on action labels.
+IL stands for imitation learning, which is also based on action primitives and UNets.
+We treat the task as a multi-class classification task (selecting a pixel in the image) using cross-entropy loss and only retain positive samples for training the UNets.
+Compared with our method, IL cannot utilize failure data for training, which results in degraded performance.
+SAC stands for Soft Actor-Critic algorithm [32].
 In SAC, the input to the value network consists of both the observation and the action.
@@
 Compared to our method, SAC requires learning an additional action network, while our method directly finds the maximum position on the output of the value network.
-”ACT” and ”DP” (diffusion policy) are methods proposed by [43] and [70].
-We directly use data generated by heuristics to train an end-to-end imitation learning policy.
-The network’s input consists of the raw depth images from the head camera as well as the masks of the clothing and the hanger.
-Our method performs best in simulation and achieves an overall 85.1% success rate.
-ACT performs excellently when inserting the first endpoint, but due to accumulated errors, it performs poorly when inserting the second endpoint.
-Similar issues are observed in other end-to-end imitation learning baselines such as diffusion policy.
-We also observe that SAC performs worst probably due to the additional action network not sufficiently trained.
+The success rate is lower because the additional action network causes additional training difficulty.
+ACT and DP (diffusion policy) are methods proposed by [33] and [34], respectively.
+We directly use data generated by heuristic policy to train an end-to-end imitation learning policy.
+The network’s input consists of raw depth images from the head camera, along with the masks of the clothing and the hanger.
+The output of the network is the angle of each robot joint.
+We only retain successful trajectories from the data for training, with a total of approximately 10,000 trajectories.
+All trajectories begin with the hanger already being grasped.
+Compared to our output of only 4 actions, these end-to-end policies need to output about 200 actions in a complete trajectory.
+These policies perform excellently when inserting the first endpoint, but due to accumulated errors and lack of data to recover from errors, they struggle with inserting the second endpoint.
 IV-C Ablation Studies in Simulation
@@
 TABLE II: Ablation study in simulation.
-”w/o online data” indicates training the network using only offline data without online data, under the same data quantity.
-Fine-tuning with online data leads to an improvement in the final success rate.
-”w/o neckline mask” indicates training the policy without using the neckline mask.
+w/o online data refers to training the network using only offline data (generated by the heuristic policy), without incorporating online data (generated by the UNet policy), while keeping the data quantity the same.
+Fine-tuning with online data results in an improvement in the final success rate.
+w/o neckline mask indicates training the policy without using the neckline mask.
 It is evident that the neckline mask also contributes to an increase in the final success rate.
 IV-D Comparison with Baselines in Real World
-In real-world experiments, we test 8 different garments and Hanger 1, as shown in Figure 4.
+In real world experiments, we test 8 different garments and Hanger 1, as shown in Figure 4.
 For a fair comparison, we use the robot to initialize the garment configuration.
 Additionally, to evaluate our policy’s performance under different initializations, we also report the policy’s success rate with human initialization.
+More qualitative results are shown in Figure 5.
+Figure 5: The figure shows the different garment states after applying each action primitive of RoboHanger system in the real world.
 Output S1(%) $\uparrow$ S2(%) $\uparrow$ S(%) $\uparrow$
@@
 TABLE III: Experiments in the real world.
-We present the performance of our policy compared to other methods on the real robot in Table III.
-We include ”IL” (imitation learning), which performed best among baseline policies in simulation; ”ACT” (a policy that directly outputs joint angles); and ”Fix” (a policy outputs hard coded parameters of both action primitives) as baselines.
-Our method outperforms all baseline methods in terms of the final success rate which demonstrates the efficacy of our method.
-Although the ACT policy shows a certain level of success in simulation, with a particularly high success rate when inserting the first endpoint, it suffers from a severe sim-to-real gap and almost never succeeds in the real world.
-Fixed policy is a manually tuned policy for the test garments.
-It can be observed that the fixed policy still shows high success rates for inserting the left endpoint because the initial position of the clothes is relatively fixed in our experiments.
-However, when inserting the right endpoint, the success rate of the fixed policy drops significantly due to unforeseen deformations of the clothes during the left endpoint insertion.
-In the last row of the table (”Ours(H)”), we further test the performance of our policy when the garment is initialized by a human.
-When the garment is initialized by the robot, the success rate shows only a slight decrease compared to human initialization, demonstrating the robustness of our policy to the initialization process.
+We present the results of our policy compared to other methods on the real robot in Table III.
+We include IL (imitation learning), which performs the best among the baseline policies in simulation; ACT (a policy that directly outputs joint angles); and Fix (a policy that outputs hardcoded parameters for both action primitives) as baselines.
+Our method outperforms all baseline methods in terms of the final success rate, which demonstrates the efficacy of our approach.
+Although the ACT policy shows a certain level of success in simulation, particularly with a high success rate when inserting the first endpoint, it suffers from a severe sim-to-real gap and rarely succeeds in the real world.
+The fixed policy is a manually tuned policy designed for the test garments.
+We find that the fixed policy still achieves high success rates for inserting the left endpoint, as the initial position of the garment remains relatively fixed in our experiments.
+However, when inserting the right endpoint, the success rate of the fixed policy drops significantly due to unforeseen deformations of clothes during insertion of the left endpoint.
+In the last row of the table (Ours(H)), we further report the performance of our policy when the garment is initialized by a human.
+When the garment is initialized by the robot, the success rate shows only a slight decrease compared to human initialization, demonstrating the robustness of our policy to variations in the initialization process.
 IV-E Generalization of Hangers in Real World
-To further study the generalizability of our method in terms of different hangers, we report the performance of our policy when using different hangers in Table IV.
+To further study the generalizability of our method across different hangers, we report the performance of our policy using various hangers in Table IV.
 Here, Hanger 2 is a hanger seen during training but without a crossbar.
-Hanger 3 and Hanger 1 have the same geometry, but due to difference settings in the 3D printing process, Hanger 3’s surface is much rougher.
+Hanger 3 and Hanger 1 have the same geometry, but Hanger 3 has a much rougher surface.
 Hanger 4 is a hanger that is not seen during training.
@@
 Although Hanger 2 was seen during training, there is a significant performance drop when inserting the right endpoint.
-As shown in Figures 5(a) and 5(b), these two images compare the state when inserting the right endpoint using different hangers.
+As shown in Figures 6(a) and 6(b), these two images compare the state when inserting the right endpoint using different hangers.
 The presence of a crossbar helps the hanger separate the upper and lower layers of the clothing, making insertion easier.
-When using Hanger 3, as shown in Figures 5(c) and 5(d), if the hanger has too much friction, it may drag the clothes while moving, causing the clothes to become very wrinkled, which in turn reduces the success rate.
+When using Hanger 3, as shown in Figures 6(c) and 6(d), if the hanger has too much friction, it may drag the clothes while moving, causing the clothes to become very wrinkled, which in turn reduces the success rate.
 When using Hanger 4, the complexity of the hanger’s shape also affects the final success rate.
 Figure 6: Results with different hangers in the real world.
-In the real world, issues may arise when changing hangers.
+Issues may arise when changing hangers in reality.
 Figure (a) and Figure (c) show results with Hanger 1.
@@
 Figure (d) illustrates that with Hanger 3, which has higher friction, the clothing tends to get wrinkled during the insertion of the left end, causing the garment to become stuck in a worse state.
-V Limitation
-For garments with less elasticity, such as shirts, more precise force feedback control might be required to avoid damage to the hanger and garment.
-Another challenge could arise when applying our action primitives to irregularly shaped garments, such as a shirt with a hood.
+V Failure Cases and Limitations
+Failure Cases.
+For the insertion of the first endpoint, the main failure mode is the inability to separate the front and back layers of the garment.
+The primary reason is the complex contact forces between the garment and the hanger.
+For inserting the second endpoint, the main failure mode is the inability to rotate the hanger into the garment neckline.
+The primary reason is the large deformation of the garment at this stage, which increases the sim-to-real gap.
+Limitations.
+Currently, garment grasping in the simulation is simplified as the garment being attached to the gripper.
+This physical modeling limits the exploration of potentially better strategies.
+A more fine-grained simulation of garment-gripper contact could further improve the overall success rate.
+For garments with less elasticity, such as shirts, more precise force feedback control may be required to avoid damaging both the hanger and the garment.
+Additionally, closed-loop control could also be introduced to help recover from failure cases.
 VI Conclusion
-In this work, we investigate the task of inserting a hanger into the neckline of clothes.
-For such a challenging but rarely studied task, we propose a simulation environment and data collection pipeline to address it.
-We demonstrate the efficiency of training the algorithm in a simulated environment, as well as its effectiveness and generalizability in the real world.
-Our pipeline facilitates efficient demonstration collecting for future robotic foundation models that require large amounts of diverse data.
-Future work may involve relaxing the assumptions for this task and extending the method to handle a wider variety of garments.
+In this work, we explore the task of inserting a hanger into the neckline of an unseen garment.
+For this challenging task, we propose a simulation environment and a data collection pipeline.
+We demonstrate the efficiency of training the algorithm in a simulation environment, as well as its effectiveness and generalizability in the real world.
+Our pipeline enables the efficient collection of demonstrations for future robotic foundation models that require large amounts of diverse data.
+Future work may involve relaxing the assumptions made for this task and extending the method to handle a broader variety of garments.
+In addition, exploring the input of RGB images could be an interesting research direction.